if the points (3, r) and (r, -1) lie on a line with slope -3/4, wat is the value of r?
Answer by nerdybill(7384)

if the points (3, r) and (r, -1) lie on a line with slope -3/4, wat is the value of r?
.
Slope between two points on a line:
m = (y2 - y1)/(x2 - x1)
.
Therefore, we have:
-3/4 = (-1 - r)/(r - 3)
Multiply both sides by 4(r-3):
-3(r-3) = 4(-1 - r)
-3r+9 = -4 - 4r
r+9 = -4
r = -13